# Scaffolding Creativity: Integrating Generative AI Tools and Real-world Experiences in Business Education

### v1

@@ 3. FINDINGS
    Students were asked to identify the raw materials and the production process of their brand merchandise to ensure that they are aligned with their brand designs. Hence all students shifted to production process at one point in their prompt histories.
  • 2
@@
    Students with prior domain knowledge (S1 and S2) exhibited non-linear learning patterns, exploring both specific and broad concepts iteratively. All three students showed accelerated acquisition of domain-specific knowledge.
3.1. AI-Assisted Research in Unfamiliar Topics
The assignment instructions included general ChatGPT usage principles:
  • encouraged open-ended exploration
  • start with broad questions to establish a foundation
  • take an iterative approach
All of the students used ChatGPT-3.5 to aid them in their research process, but only 3 out of the 8 students documented their entire prompt histories in their submissions, all of which exhibited adaptive prompting skills.
Analysis of the documented prompt histories and final presentations revealed several interesting patterns:
Non-Linear Learning Trajectories
Despite the assignment instructions being structured in a typical linear format – from broad to narrow – to help students narrow down their research focus, the sequencing of the questions some students asked revealed non-linear learning trajectories. This divergence from the expected pattern suggests that AI-assisted research might fundamentally alter how students navigate unfamiliar topics.
Students with strong domain knowledge in certain fields demonstrated distinct starting points in their AI interactions. For instance, S1, leveraging prior internship experience in garment production, began with questions of the raw materials she already knew were eco-friendly, by asking what garments can be made from these materials. This approach indicates an attempt to validate and expand existing knowledge using ChatGPT as a reference point.
In contrast, S2 drawing on design experience, initiated the interaction by describing her product vision with clear design-oriented specifications such as “water bottle pockets” and “enough space for students”. Possibly due to the lack of experience with ChatGPT, S2 did not ask any question after specifying her product requirement. Realizing that ChatGPT’s response was not what she was looking for, she returned to a more general question of what materials are usually used for her product categories.
S3, who had limited experience in either production or design, followed a more general and linear questioning pattern in the prompt history. This pattern suggests that students without prior domain knowledge may rely more heavily on the structured approach provided in the instructions.
Students’ background in domain knowledge also seemingly affected their analytical approach when interacting with ChatGPT. Students with less domain knowledge tend to accept AI responses more readily, using them as scaffolding to build general understanding. But the students with certain domain knowledge are more likely to demonstrated a more critical and nuanced engagement with AI-generated information.
For example, when S1 learned that hemp can be used to create “even jeans” she immediately followed up with a question about the differences between cotton and hemp jeans for deeper inquiry.
Similarly, S2 cross-referenced AI responses with internet sources, particularly when ChatGPT’s answers seemed too theoretical for her practical, production-oriented questions. This triangulation between AI, web resources, and personal knowledge illustrates a sophisticated approach to information validation that leverages domain expertise.
the internet shows that nylon and polyster blend fabrics are recommended, what do you think? [S2]
Moreover, one of the students demonstrated a tendency to inquire about tangential knowledge. In final queries, S1 focused on ramie, a fabric not mentioned in previous responses. This sudden shift in focus could represent either a recollection of prior knowledge or genuine curiosity sparked by the AI interaction. This behavior could suggest the AI’s responses might trigger memories of related concepts or introduce new ideas that pique the student’s interest, leading to unexpected learning paths.
Dynamic Shift of Scope of Questioning
Throughout their interactions with ChatGPT, students demonstrated a notable pattern of oscillating between broad and narrow inquiries. This dynamic shift in the scope of questioning reveals a unique aspect of AI-assisted learning.
For instance, S1 began with a specific question about eco-friendly fabrics, then broadened her inquiry to ask about ”the most eco-friendly fibers for clothing,” before narrowing again to inquire about specific properties of organic cotton, hemp, and Tencel.
Similarly, S2 started with a broad product description, narrowed to specific material inquiries, then shifted the focus to general production processes, which is the second part of the assignment.
S3 added her own comments in her prompt history documentation, mentioned she “had no idea there were a variety of ways to produce hoodie’s graphic designs” after P2. ChatGPT’s response, presented as a long list of bullet points, initially led her to incorrectly assume that some of the materials are used the same processes. Upon realizing her misunderstanding from follow-up responses, she broadened her inquiry to ask about the general production process of hoodies.
This oscillation between broad and narrow questioning reflects an adaptive learning which allows the students to gain contextual understanding and specification of details. There is also seemingly redirection. Students also tend to redirect their focus or scope after getting responses with unexpected information or concepts.
These cases may suggest that AI-assisted learning could allow students to rapidly shift between scopes to facilitate a more flexible and responsive research process. However, this pattern also presents challenges. The ease of shifting scope might lead to a scattered learning process if not managed well.
Accelerated Domain-Specific Knowledge Acquisition
This observation opens up exciting possibilities for personalized, project-based learning, where students can quickly acquire the specific expertise needed for their individual projects within compressed timelines. It also highlights the potential of AI as a powerful tool for just-in-time learning in professional contexts.
While it’s understandable that business students researching unfamiliar disciplines aren’t necessarily required to understand all the details, this rapid acquisition of knowledge still raises important considerations:
1. What role does critical evaluation play in this accelerated process?
2. How does this rapid learning affect knowledge retention?
3. How can we contextualize the tradeoff between the speed of knowledge acquisition and the development of deeper understanding?
These questions highlight the need for further research into the cognitive processes involved in AI-assisted learning and the development of pedagogical strategies.
3.2. Dynamic Interaction with Real-World Experiences
To investigate how AI can serve as more than just an information source and understand how human-AI collaboration works in learning contexts, we designed the course specifically to bridge the gap between the AI insights and practical application in ways previously unexplored in traditional business education.
Expanding AI Capabilities in Experiential Learning
After their AI-assisted research into raw materials and production processes, students successfully identified potential manufacturers on 1688.com, a major B2B e-commerce platform. This demonstrates how AI can expand human capabilities in real-world business scenarios, allowing novices to engage in complex market research tasks typically reserved for experienced professionals.
Uncovering AI Limitations through Real-World Engagement
To further uncover the practicality of the AI-generated knowledge and synethsize it into the entire experiential learning, the students went to Yiwu Small Merchandise Market to gain firsthand market intelligence, and have a chance to have a real-world validation in AI-assisted learning.
Not surprisingly, not everything what ChatGPT provided proved to be useful in reality. For example, S1’s discovery that 100% hemp production was far less common than AI had implied. Moreover, many students encountered a wide array of materials and production techniques during their visit that were not mentioned in their AI-assisted research.
This discovery points to two known limitations of AI-generated knowledge: first, its tendency to provide generic answers unless prompted with very specific questions, which is difficult for student users; and second, its potential lag in incorporating the latest innovations or highly specialized industry knowledge.
It suggests a need for AI interfaces that encourages a continuous learning loop, where real-world experiences and additional knowledge sources not only supplemented AI-generated information but also informed future AI interactions.
Human-AI Synergy: Navigating Complex Business Environments
Perhaps the most significant insight emerged from the unexpected ways students combined AI-generated knowledge with real-world experiences to navigate the opaque Yiwu market. Contrary to expectations, the rapid knowledge acquisition facilitated by AI allowed some students to successfully and confidently pose as experienced buyers during vendor interactions, gaining valuable market intelligence.
This phenomenon was particularly notable given that the vendors can easily identify novice business buyers or ”tourists” (retail buyers) if they don’t ask the ”right” questions, and offer unfavorable deals. The AI-assisted preparation appeared to provide students with sufficient domain knowledge to ask pertinent questions and engage in more nuanced conversations, potentially altering vendor perceptions and interactions.
The observed synergy between AI preparation and human adaptability suggests a novel model of human-AI collaboration in learning contexts. Students utilized AI-generated knowledge as a foundational framework, which they then dynamically adjusted based on real-time observations and interactions in experiential learning. This process of continuous recalibration between AI inputs and human experiences suggests a more fluid and interactive model of AI-assisted learning than previously conceptualized.
The opacity of the Yiwu market, with its unwritten rules and subtle negotiation cues, highlighted areas where human experiential learning still surpasses AI capabilities. However, the AI-assisted preparation enabled students to engage more confidently with these tacit aspects of business interactions, suggesting that AI can play a crucial role in preparing learners for complex, real-world scenarios.
3.3. AI-Assisted Visual Transformation
Traditionally, business school assignments have predominantly relied on text-based presentations supplemented with basic visuals. This approach, partly attributed to students’ non-design backgrounds, has limited the exploration of visual communication in business education. Our case study investigates how the introduction of AI-assisted text-to-image and image-to-video tools impacts students’ ability to visualize and communicate business concepts.
Four students (N=4) kept record of their interaction with Midjourney in the class Discord channel after the training session. Other participants utilized various AI image generation tools through personal accounts, such as Leonardo.ai, which precluded consistent data collection across the entire cohort. Table 4 presents the statistics of these interactions, including initial concept prompts, iterative prompts (usually style or component changes), upscaling actions, remixes, uploads, image-to-text descriptions and command error corrections.
Table 4. Midjourney Interactions (Post-Training)
ID S2 S3 S4 S5
Interactions 90 31 40 12
Concepts 14 10 12 4
Iterations 45 10 19 6
Effective Interactions per Concept 2.6 1.0 1.6 1.5
  • 1
    Students could make errors in their prompts due to their unfamiliarity with the command format in Midjourney’s server on Discord, causing ineffective interactions.
AI as a Cognitive Scaffold and Human-AI Co-Creation
Our study revealed that AI tools functioned not merely as implementation aids, but as collaborative partners in both concept development and visualization. This partnership manifested as an iterative process where AI-generated visuals influenced conceptual thinking, which in turn guided further visual exploration.
In particular, the introduction of multimodal AI tools after midterm appeared to serve as cognitive scaffolds, enabling students to bridge the gap between abstract business concepts and concrete visual representations. This scaffolding effect was evident in the expansion and refinement of initial ideas:
Moreover, analysis of prompt histories revealed that students often incorporated unexpected elements from AI generations into their subsequent prompts. Initially focusing on product features in her prompts, S2 noticed a blurry figure in one AI-generated image. This unexpected element inspired her to incorporate human figures in subsequent prompts, which are better representation of her easy-to-use brand concepts.
Analysis of S4’s prompt history revealed significant variations in length and complexity ($\mu=51.23$, $\sigma=30.74$). This variability suggests potential utilization of text-based AI tools or prompt aid tools to generate Midjourney prompts. This observation indicates that even novice users are capable of leveraging multiple AI tools in tandem to enhance their creative process.
This collaborative process aligns with the vision that AI tools act not just as implementers of human ideas, but as active collaborators in the creative process. It also resonates with the notion of AI as a coach or a colleague (Lubart, 2005), stimulating new directions of thought and encouraging students to reconsider and enhance their initial concepts.
Democratization of Visual Design Skills
The rapid adaptation to visual communication also suggests that AI tools may be democratizing visual design skills in non-design disciplines. All students demonstrated a strong capability to shift towards visual communication, regardless of their prior design experience.
This democratization was particularly evident in the professional-quality outputs produced by the students. The quality and sophistication of these visual materials markedly exceeded typical outputs observed in traditional business education courses, particularly among first- and second-year students.
As noted previously, none of the students had prior experience with text-to-image or image-to-video tools in an academic or professional context before the in-class training session in Week 9. While some students may have casually experimented with these tools for personal amusement, they had not applied them to serious project work before this course. This rapid skill acquisition in a professional context suggests that students may possess an innate capacity for visual expression that was previously untapped in traditional business education. AI tools appear to be unlocking this potential, raising important questions about the future of visual literacy in business education.
Shift Towards Individual Expression and Engagement
The introduction of AI tools in the course also led to an unexpected shift in student behavior, challenging traditional collaborative norms in business education. Despite encouragement to form pairs after the brand proposal stage, 6 out of 8 students insisted on working individually on their projects. This preference for individual work, especially given the comprehensive nature of creating an entire brand portfolio, marks a significant departure from typical business education practices.
This shift towards individual work aligns with the concept of ”AI as an Empowering Partner” (UNESCO, 2023). The AI tools appeared to increase students’ sense of agency and self-efficacy, enabling them to tackle complex design tasks that would typically require team effort or specialized skills.
The high motivation observed in students to express and implement their ideas independently may also resonates with Csikszentmihalyi’s Flow Theory(Csikszentmihalyi and LeFevre, 1989). The AI tools seem to have created an optimal challenge-skill balance, facilitating a state of flow where students felt both capable and motivated to pursue ambitious individual projects.
Moreover, while traditional business education often relies on human-human collaboration to compensate for individual skill gaps, our study suggests that AI tools are redefining this dynamic. This phenomenon can be understood through the lens of Human-AI Collaboration [16]. This also paves the way for a new pedagogical shift in personalized learning.
4. IMPLICATIONS AND LIMITATIONS
4.1. Implications for Business Education
Our case study reveals a fundamental shift in the landscape of business education, one that demands a radical rethinking of current pedagogical models. The integration of AI tools, coupled with real-world experiences, points towards a future where business education transcends traditional disciplinary boundaries, becoming inherently interdisciplinary and deeply intertwined with technological literacy.
The observed non-linear learning trajectories and rapid skill acquisition facilitated by AI tools suggest that the future of business education lies not in the traditional transmission of static knowledge, but in the development of adaptive learning capabilities. This shift addresses the ’curriculum chaos’ noted by [10] by embracing flexibility as a core principle. Rather than struggling to keep pace with rapidly changing business landscapes, educational institutions should focus on creating dynamic learning environments where AI serves as a collaborative partner in the educational process. This approach could potentially resolve the tension between the need for current, relevant knowledge and the practical constraints of curriculum development.
Furthermore, the democratization of design and technical skills through AI tools points to a future where the boundaries between business, design, and technology become increasingly blurred. This convergence calls for a reimagining of business education as a holistic, integrative discipline. Future business curricula should not merely incorporate AI as a tool, but should be fundamentally restructured around the principle of human-AI collaboration. This could involve the development of new hybrid courses that blend elements of business strategy, design thinking, and AI interaction, preparing students for a world where these elements are inseparable in practice.
The shift towards individual expression enabled by AI tools, coupled with the continued importance of real-world collaborative experiences, suggests a need for new models of collaborative learning. Future business education should explore AI-mediated collaboration platforms that can balance individual creativity with team-based problem-solving. This could potentially address the challenge of developing collaborative skills in an AI-augmented world, where individual capabilities are significantly enhanced by technology.
Perhaps most critically, our findings underscore the need for a new approach to developing business acumen in the AI era. Rather than focusing on the acquisition of specific skills or knowledge sets, which may quickly become obsolete, business education should prioritize the development of ’AI-interaction literacy’. This involves not just technical proficiency with AI tools, but a deep understanding of how to critically evaluate AI-generated insights, how to effectively prompt and guide AI systems, and how to synthesize AI-produced information with human expertise and real-world observations.
The integration of real-world experiences with AI-assisted learning in our study also points to a future where the boundaries between classroom and field blur. Business schools should consider adopting more fluid educational models where AI-enhanced project-based learning is seamlessly integrated with real-world business experiences. This could involve AI-guided virtual internships, augmented reality business simulations, or AI-facilitated global collaboration projects.
4.2. Limitations and Challenges
First, the small sample size and single course context limit the generalizability of our results. Future studies should expand to larger cohorts across multiple courses and institutions to validate and extend our findings. This expansion would also help address potential self-selection bias, as students who chose to take this innovative course might be more receptive to AI integration than the general student population.
The lack of longitudinal data is another limitation. While we observed rapid skill acquisition and knowledge development, the long-term retention and application of these skills remain unexplored.
The potential novelty effect of AI tools is also a consideration. The heightened engagement and outcomes we observed might partially result from the novelty of the technology. Longitudinal studies are needed to determine if these positive effects persist as AI tools become more commonplace in educational settings.
Assessment challenges in an AI-assisted learning environment emerged as a significant limitation. Distinguishing between AI-generated content and student-generated insights proved difficult, raising questions about how to fairly evaluate learning outcomes. This challenge calls for the development of new assessment methodologies tailored to AI-integrated curricula, an important area for future research.
The scalability of our approach presents challenges beyond mere technical implementation. A striking example emerged when a student, aiming to generate images representing eco-friendly garment production, consistently received AI-generated visuals of Asian factory settings with Asian workers. This outcome, while reflective of current global manufacturing trends, inadvertently highlighted the intricate interplay between AI biases, global economic realities, and brand representation.
This highlights the necessity to develop critical thinking skills that transcend traditional business silos, requiring both educators and students to analyze AI-generated insights within broader social, ethical, and economic contexts.
In conclusion, the future of business education should be dynamic, AI-enhanced learning ecosystems that prioritize adaptability, interdisciplinary integration, and the development of new forms of literacy. This vision challenges institutions to move beyond incremental changes and towards a fundamental reimagining of what business education can and should be in an AI-augmented world. For the HCI community, this opens up new avenues of research into the design of AI systems that can support this educational transformation, from adaptive learning interfaces to AI-mediated collaboration tools.

### v2

@@ Appendix A Student Prompt Histories and Observed Behaviors
1 Students were asked to identify the raw materials and the production process of their brand merchandise to ensure that they are aligned with their brand designs. Hence all students shifted to production process at one point in their prompt histories.
@@
2 Students with prior domain knowledge (S1 and S2) exhibited non-linear learning patterns, exploring both specific and broad concepts iteratively. All three students showed accelerated acquisition of domain-specific knowledge.
Appendix B Mini-Presentation 4 Assignment Instructions
Overview: The objective of this assignment is to conduct a comprehensive research project on one of the suppliers or factories listed on 1688.com. For those not proficient in Chinese, the website can be translated into English when accessed via a computer. As an alternative, you may use Alibaba.com, which offers language selection at the top right corner of the homepage. Please note that suppliers on Alibaba.com may have a limited product range and higher prices due to their international shipping capabilities.
Your Presentation should include:
  1. (1)
    A very short introduction into a product category you are interested in: this can be a popular consumer good in your home country; or a product you would like to introduce for your own brand or for a certain influencer (for example, a custom-designed water bottle for the gym goers); or simply something you are curious about. (suggested duration 1min)
  2. (2)
    2. Basic research into the raw materials: list what types of raw materials you need for the product. Make it as specific as possible. For example, instead of just mentioning “plastic”, you should identify the exact type such as PVC (Polyvinyl Chloride) or Nylon. If you want to maintain high safety standards for the product, you need to be careful with what materials you use. Use pictures as well to illustrate the materials. (suggested duration 1min)
  3. (3)
    Basic research into the production process: what types of special production processes are involved in creating the product. Detail the methods used, such as sewing buttons onto jeans either by hand or machine, stringing beads into a necklace by hand. Use images to illustrate the production process or craftsman ship. (suggested duration 1min)
  4. (4)
    Find a few suppliers/factories on 1688.com (or alibaba.com) who can provide such materials and production service. Sometimes, you source different materials from different suppliers and get them assembled by another factory (if cheaper). You can also choose to buy kits with higher prices from a particular supplier. Tell us why you believe such factories are good choices, for example, low prices, authenticity (evidenced by factory photos), many followers, contracts with other big-name brands, long years of service, experience with international e-commerce. (suggested 2min).
Tips:
  • Your initial research for Questions 2 and 3 DOES NOT need to be 100% accurate. The primary goal is for you to engage with new and unfamiliar topics, so the emphasis will be on your research effort rather than precision. This approach will be reflected in the grading rubric.
  • I highly recommend you use ChatGPT to you get started with Question 2 and 3. Remember, it is very rare to get the perfect answers with the first attempt. Start with broad questions to establish a foundation, such as, ”What types of plastics are used for making packaging?” From the general answers you receive, refine your questions to dig deeper. For instance, you could follow up with, ”What are the different types of PVC, and what are their common applications, advantages, and disadvantages?” You should practice your prompting skills through this exercise.
  • Your presentation time should not exceed 5min. I will put a timer by the desk in the next class. You cannot continue beyond the 5min cutoff.
Grading Rubrics (weighed equally)
  • Detail and Specificity - The research provides specific details about the raw materials, the production processes, avoiding generic terms.
  • Visual Illustration with engaging pictures
  • Safety (or ESG) considerations
  • Research depth - Evidence of thorough research in finding suitable suppliers or factories with clear justification for choices.
  • Presentation Organization, Delivery and Time Management – well structured, logically organized, confident delivery within the time limit.